# Potato Potahto in the FAO-GAEZ Productivity Measures?

### v2

@@ Abstract
The FAO-GAEZ crop productivity data are widely used in Economics. However, the existence of measurement error is rarely recognized in the empirical literature. We propose a novel method to partially identify the effect of agricultural productivity, deriving bounds that allow for nonclassical measurement error by leveraging two proxies. These bounds exhaust all the information contained in the first two moments of the data. We reevaluate three influential studies, documenting that measurement error matters and that the impact of agricultural productivity on economic outcomes may be smaller than previously reported. Our methodology has broad applications in empirical research involving mismeasured variables.
@@ Introduction
To address this issue, we propose a novel econometric method that partially identifies the effect of agricultural productivity while accounting for this variable’s mismeasurement. To empirically ground our methodology, we assess three empirical studies (Nunn and Qian, 2011; Bustos et al., 2016; Acharya et al., 2016) and find the possibility of nonclassical measurement error in the FAO-GAEZ productivity measures by comparing its last two versions (Fischer et al., 2012, 2021). Lastly, we leverage the information contained in both proxies through our proposed statistical tool and find that the effect of agricultural productivity on economic outcomes might be smaller than previously considered in these three studies.
@@
Lastly, we highlight that our proposed methodology can be used in many empirical contexts (Appendix F). To employ our partial identification strategy, we only require an exogenous mismeasured treatment variable and two exogenous proxies. As examples of possible applications, we mention the New Keynesian Phillips Curve (Gali and Gertler, 1999) and the impact of media access on economic outcomes (La Ferrara et al., 2012; Adena et al., 2015). In the first case, aggregate marginal cost is not directly measured, requiring the use of model-based proxies. In the second case, media access is measured through an irregular terrain model, which has been modified over time.
@@
Related Literature. This article contributes to two different branches of literature. Concerning its empirical contribution, many papers in economics have used the FAO-GAEZ data. Nunn and Qian (2011), in pioneering work, uses the potential productivity of potatoes to explain how its introduction in the Old World affected population and urbanization. Bustos et al. (2016); Bustos et al. (2019); Bustos et al. (2020) use the productivity of soybeans to measure the impact of introducing genetically modified seeds on structural transformation. Bentzen et al. (2017) uses productivity under rain-fed and irrigation conditions to study the impacts of irrigation on the origins of autocracies. Acharya et al. (2016) uses cotton productivity as an instrument for slavery to explain contemporary differences in political attitudes in the U.S. Carillo (2021) uses wheat productivity to study the impact of agricultural policies on industrialization in Italy. [^3] Similar papers include Moorthy (2022), Dias et al. (2023) and Carreira et al. (2024) on soybean productivity and gender disparity, water pollution, and deforestation; Galor and Özak (2016) on the productivity of crops available before and after the Columbian Exchange and time preferences; Alesina et al. (2011, 2013) on the productivity of various crops, plough use, and gender disparities; Dube et al. (2016) on maize productivity and the Mexican drug sector; Gollin et al. (2021) on the productivity of various crops and the impacts of the Green Revolution; Mayshar et al. (2022) on the relative productivity of cereals over roots and tubers and the origin of states; Fiszbein (2022) on the productivity of multiple crops and its effects on development. These papers use FAO-GAEZ data to estimate linear specifications of the impacts of crop productivity on various outcomes of interest. However, it is worth noting that there is a range of papers employing FAO-GAEZ data in different types of models, such as trade models (Costinot et al., 2016; Sotelo, 2020; Pellegrina, 2022; Farrokhi and Pellegrina, 2023), deforestation models (Sant’Anna, 2024; Araujo et al., 2025; Dominguez-Iino, 2021), and other types of spatial models (Adamopoulos and Restuccia, 2022; Adamopoulos et al., 2022; Imbert et al., 2022; Fernández-Villaverde et al., 2023).
@@
We also highlight that our work is connected to the studies developed by Black et al. (2000), Cunha et al. (2010), Alix-García and Millimet (2023) and Millimet and Parmeter (2024). Similarly to us, Black et al. (2000) and Millimet and Parmeter (2024) focus on nonclassical measurement error problems. However, differently from our methodology, they impose that the true variable of interest and the measurement error are negatively correlated with each other. Furthermore, Alix-García and Millimet (2023) focus on mismeasured binary outcome variables while our model analyzes a mismeasured independent variable. Lastly, Cunha et al. (2010) require three proxies to achieve point-identification in a model with nonclassical measurement error, while our approach requires only two proxies to partially identify the coefficient of interest.
Paper Organization. The remainder of this paper is organized as follows. Section 2 describes the FAO-GAEZ data, documents the existence of measurement error in its variables, and intuitively explains the use of partial identification as a solution to measurement error. Section 3 presents our econometric model and states our main identification results. Section 4 briefly describes the estimation and inference methods connected to our partial identification strategy. Section 5 illustrates the proposed tools by reevaluating the impact of agricultural productivity on three empirical contexts. Lastly, Section 6 concludes.
@@
There is no precise documentation of the differences between versions. Upon inspection of versions 3 and 4, we identified many differences in parameters and raw data. To illustrate this issue, we provide four examples. First, the “harvest index” parameter essentially measures the portion of the weight of the crop that is actually of economic interest — e.g., the beans of soybeans — and is different for a set of crops between versions. Second, for another set of crops, FAO-GAEZ updated the “leaf area index”, which measures the plant’s canopy and its ability to capture sunlight. Third, precipitation data in version 3 is from the Variability Analysis of Surface Climate Observations 1.1 (VASClimO), while version 4’s precipitation data is from the Global Precipitation Climatology Centre (GPCC) Full Data Reanalysis Product Version 6. Fourth, data that still uses the same source are using newer versions. For example, temperature, sunshine fraction, and wind speed were from CRU 1.1 in version 3 and from CRU 3.21 in version 4, which entails a range of internal modifications of each data source. [^5] Rhode (2024) discusses differences in the model for the productivity of cotton.
@@
Nunn and Qian (2011) studies the impact of the introduction of the potato in the Old World on population, using the FAO-GAEZ data on suitability of white potato. Figure 1 shows the spatial differences between versions 3 and 4 for potato productivity. Some countries — such as Norway and Iraq — had little to no area suitable for potato in version 3 but a large suitable area in version 4, creating a large change in logs. Another example is China, which has an increase in area suitable for potatoes of 700% across versions.
@@
This figure shows the spatial distribution of the differences in FAO-GAEZ versions 3 and 4 for the log of the area suitable for potato (low input) on the left. The plot on the right shows the correlation between versions 3 and 4 for the log of area suitable for potato. Notice we follow Nunn and Qian (2011) in using the transformation $\log(1+x)$ since our objective is that our results are comparable with the ones in the original paper.
@@
This figure shows the spatial distribution of the differences in versions 3 and 4 for the change in soybean productivity (high input minus low input) on the left. Each color represents a decile of the data distribution. The plot on the right shows the correlation between versions 3 and 4 for the change in soybean productivity.
@@
This figure shows the spatial distribution of the differences in versions 3 and 4 for the cotton productivity high input on the left. Each color represents a quintile of the data distribution. The plot on the right shows the correlation between versions 3 and 4 for cotton productivity.
@@
where $Z_{1,res}=M\cdot Z_{1}$ and $Z_{2,res}=M\cdot Z_{2}$. Intuitively, $Z_{1,res}$ and $Z_{2,res}$ are two imperfect measures of $X_{res}^{*}$ where $U_{1}$ and $U_{2}$ capture the measurement error.
@@
Assumption 3 imposes that the effect of the true variable of interest on the outcome variable is non-negative. [^7] The case that imposes that the true target parameter is non-positive is symmetric. Moreover, if the researcher is unwilling to impose Assumption 3, Appendix D offers an alternative partial identification strategy that does not require prior knowledge of the direction of the true effect. As a consequence of imposing fewer assumptions, the identified interval around the parameter of interest is weakly wider. In our empirical application, it imposes that the effect of the true crop productivity on the outcome variable is non-negative. This assumption is motivated by the empirical literature because we assume that the articles using one proxy — the FAO-GAEZ version 3 — are at least identifying the correct sign of the true effect. If a researcher is not willing to impose this type of assumption, Appendix D proposes an alternative way to partially identify this coefficient without making assumptions about its sign.
@@
Assumption 4 imposes that the measurement error terms in our proxies of the true variable of interest are uncorrelated with the unobservable economic shock after controlling for observable covariates. [^8] This assumptions is implied by $Cov\left(A,\epsilon\right)=0$ for any $A\in\left\{W_{1},\ldots,W_{J},X^{*},Z_{1},Z_{2}\right\}$ as discussed in Appendix H. This assumption weakens the implicit identifying assumption in the empirical literature about agricultural productivity (Nunn and Qian, 2011; Bustos et al., 2016; Acharya et al., 2016) since applied researchers impose that $U_{1}=U_{2}=0$. In our empirical application, it intuitively imposes that prediction errors in the FAO-GAEZ measures are uncorrelated with the unobserved economic forces that determine the outcome variable.
@@
To illustrate the use of our partial identification strategy, we reevaluate the findings of Nunn and Qian (2011), Bustos et al. (2016), and Acharya et al. (2016). Although we find that results by Bustos et al. (2016) are robust to nonclassical measurement error, our method suggests that the impact of agricultural productivity on economic outcomes may be smaller than argued by Nunn and Qian (2011) and Acharya et al. (2016).
@@
Nunn and Qian (2011) studies the impact of the introduction of the potato in the Old World on population. The empirical specification uses panel data to estimate Equation (8):
@@
Column (6) shows the results associated with our method. As we develop a partial identification strategy, our result is an estimated interval. This interval shows a wide range of possible effects, with an increase of one standard deviation in $\log(1+A_{j}^{potato})$ leading to an increase in population between 2.9% and 22.9%. Importantly, this interval exhausts all the information contained in the first and second moments of the data-generating process. Consequently, the large estimate for the upper bound is not a result of our estimator being loose.
@@
Column (3) shows the results associated with our method. The estimated interval accounts for nonclassical measurement error in the explanatory variable and shows a wide range of possible effects. Interestingly, the estimated lower bound is almost a third of the point estimate associated with FAO-GAEZ version 3 (Column (1)) and is even lower than the point estimate associated with FAO-GAEZ version 4 (Column (4)), suggesting that accounting for measurement error in this empirical context is relevant. Lastly, when we account for sample uncertainty when estimating our bounds in Column (3), we do not reject the null of no effect at the 10%-confidence level.

### v3

@@ Abstract
The FAO-GAEZ productivity data are widely used in Economics. However, the empirical literature rarely discusses measurement error. We use two proxies to derive novel bounds around the effect of agricultural productivity in a setting with nonclassical measurement error. These bounds rely on assumptions that are weaker than the ones imposed in empirical studies and exhaust the information contained in the first two moments of the data. We reevaluate three influential studies, documenting that measurement error matters and that the impact of agricultural productivity may be smaller than previously reported. Our methodology has broad applications in empirical research involving mismeasured variables.
@@ Introduction
To address this issue, we propose a novel econometric method that uses multiple proxies to partially identify the effect of agricultural productivity, accounting for measurement error. To illustrate the empirical relevance of our approach, we revisit three influential studies: Nunn and Qian (2011), Bustos et al. (2016), and Acharya et al. (2016). Applying our method with two available proxies, we find that the estimated effect of agricultural productivity on population might be substantially smaller than reported by Nunn and Qian (2011); that there might be no effect on political behavior, in contrast to the findings of Acharya et al. (2016); and that the effect on structural transformation remains robust, consistent with Bustos et al. (2016).
@@
Lastly, we highlight that our proposed methodology can be used in many empirical contexts. To employ our partial identification strategy, we only require an exogenous mismeasured treatment variable and two exogenous proxies. As examples of possible applications, we mention the impact of media access on economic outcomes (Olken, 2009; Adena et al., 2015) and the effect of subjective beliefs about life expectancy on economic behavior (Post and Hanewald, 2013; Gan et al., 2015; Bissonnette et al., 2017). In the first case, media access is measured through an irregular terrain model, which has been modified over time. In the second case, different imputation methods are used to address the focal answer problem when trying to measure subjective probabilities. Appendix F explains six other examples in the fields of Macroeconomics, Behavioral Economics, Economic History, and Political Economy.
@@
Related Literature. This article contributes to two different branches of literature. Concerning its empirical contribution, many papers in economics have used the FAO-GAEZ data. Nunn and Qian (2011), in pioneering work, uses the potential productivity of potatoes to explain how its introduction in the Old World affected population and urbanization. Bustos et al. (2016, 2019, 2020) use the productivity of soybeans to measure the impact of introducing genetically modified seeds on structural transformation. Bentzen et al. (2017) uses productivity under rain-fed and irrigation conditions to study the impacts of irrigation on the origins of autocracies. Acharya et al. (2016) uses cotton productivity as an instrument for slavery to explain contemporary differences in political attitudes in the U.S. Carillo (2021) uses wheat productivity to study the impact of agricultural policies on industrialization in Italy. [^3] Similar papers include Moorthy (2022), Dias et al. (2023) and Carreira et al. (2024) on soybean productivity and gender disparity, water pollution, and deforestation; Galor and Özak (2016) on the productivity of crops available before and after the Columbian Exchange and time preferences; Alesina et al. (2011, 2013) on the productivity of various crops, plough use, and gender disparities; Dube et al. (2016) on maize productivity and the Mexican drug sector; Gollin et al. (2021) on the productivity of various crops and the impacts of the Green Revolution; Mayshar et al. (2022) on the relative productivity of cereals over roots and tubers and the origin of states; Fiszbein (2022) on the productivity of multiple crops and its effects on development. These papers use FAO-GAEZ data to estimate linear specifications of the impacts of crop productivity on various outcomes of interest. However, it is worth noting that there is a range of papers employing FAO-GAEZ data in different types of models, such as trade models (Costinot et al., 2016; Sotelo, 2020; Pellegrina, 2022; Farrokhi and Pellegrina, 2023), deforestation models (Sant’Anna, 2024; Araujo et al., 2025; Dominguez-Iino, 2021), and other types of spatial models (Adamopoulos and Restuccia, 2022; Adamopoulos et al., 2022; Imbert et al., 2022; Fernández-Villaverde et al., 2023).
@@
We also highlight that our work is connected to the studies developed by Hausman et al. (1998), Black et al. (2000), Cunha et al. (2010), Alix-García and Millimet (2023) and Millimet and Parmeter (2024). Similarly to us, Black et al. (2000) and Millimet and Parmeter (2024) focus on nonclassical measurement error problems. However, differently from our methodology, they impose that the true variable of interest and the measurement error are negatively correlated with each other. Furthermore, Hausman et al. (1998) and Alix-García and Millimet (2023) focus on mismeasured binary outcome variables while our model analyzes a mismeasured independent variable. Lastly, Cunha et al. (2010) require three proxies to achieve point-identification in a model with nonclassical measurement error, while our approach requires only two proxies to partially identify the coefficient of interest.
Paper Organization. The remainder of this paper is organized as follows. Section 2 describes the FAO-GAEZ data, documents the existence of measurement error in its variables, and intuitively explains the use of partial identification as a solution to measurement error. Section 3 presents our econometric model and states our main identification results. Section 4 briefly describes the estimation and inference methods connected to our partial identification strategy. Section 5 illustrates the proposed tools by reevaluating the impact of agricultural productivity on three empirical contexts, and Section 6 explains how to use our tools in a practical way in other applications. Lastly, Section 7 concludes.
@@
There is no precise documentation of the differences between versions. Upon inspection of versions 3 and 4, we identified many differences in parameters and raw data. To illustrate this issue, we provide four examples. First, the “harvest index” parameter essentially measures the portion of the weight of the crop that is actually of economic interest — e.g., the beans of soybeans — and is different for a set of crops between versions. Second, for another set of crops, FAO-GAEZ updated the “leaf area index”, which measures the plant’s canopy and its ability to capture sunlight. Third, precipitation data in version 3 is from the Variability Analysis of Surface Climate Observations 1.1 (VASClimO), while version 4’s precipitation data is from the Global Precipitation Climatology Centre (GPCC) Full Data Reanalysis Product Version 6. Fourth, data that still uses the same source are using newer versions. For example, temperature, sunshine fraction, and wind speed were from CRU 1.1 in version 3 and from CRU 3.21 in version 4, which entails a range of internal modifications of each data source. [^5] Rhode (2024) discusses differences in the model for the productivity of cotton. Despite these differences, all versions of the FAO-GAEZ intend to measure the same potential productivity, i.e., the agricultural productivity for a given place, climate, technology, and water supply.
@@
Nunn and Qian (2011) study the impact of the introduction of the potato in the Old World on population, using the FAO-GAEZ data on suitability of white potato. Figure 1 shows the spatial differences between versions 3 and 4 for potato productivity. Some countries — such as Norway and Iraq — had little to no area suitable for potato in version 3 but a large suitable area in version 4, creating a large change in logs. Another example is China, which has an increase in area suitable for potatoes of 700% across versions.
@@
This figure shows the spatial distribution of the differences in FAO-GAEZ versions 3 and 4 for the log of the area suitable for potato (low input) on the left. The plot on the right shows the correlation between versions 3 and 4 for the log of area suitable for potato $(A^{potato})$. Notice we follow Nunn and Qian (2011) in using the transformation $\log(1+x)$ since our objective is that our results are comparable with the ones in the original paper. The 45 line is shown in red.
@@
This figure shows the spatial distribution of the differences in versions 3 and 4 for the change in soybean productivity (high input minus low input) on the left. Each color represents a decile of the data distribution. The plot on the right shows the correlation between versions 3 and 4 for the change in soybean productivity $(\Delta A^{soy})$. The 45 line is shown in red.
@@
This figure shows the spatial distribution of the differences in versions 3 and 4 for the cotton productivity high input on the left. Each color represents a quantile of the data distribution. The plot on the right shows the correlation between versions 3 and 4 for cotton productivity $(A^{cotton})$. The 45 line is shown in red.
@@
where $Z_{1,res}=Z_{1}-W\cdot\left(\mathbb{E}\left[W^{T}\cdot W\right]\right)^{-1}\cdot\mathbb{E}\left[W^{T}\cdot Z_{1}\right]$ and $Z_{2,res}=Z_{2}-W\cdot\left(\mathbb{E}\left[W^{T}\cdot W\right]\right)^{-1}\cdot\mathbb{E}\left[W^{T}\cdot Z_{2}\right]$. Intuitively, $Z_{1,res}$ and $Z_{2,res}$ are two imperfect measures of $X_{res}^{*}$ where $U_{1}$ and $U_{2}$ capture the measurement error.
@@
Assumption 3 imposes that the effect of the true variable of interest on the outcome variable is non-negative. [^7] The case that imposes that the true target parameter is non-positive is symmetric. Moreover, if the researcher is unwilling to impose Assumption 3, Appendix D offers an alternative partial identification strategy that does not require prior knowledge of the direction of the true effect. As a consequence of imposing fewer assumptions, the identified interval around the parameter of interest is weakly wider. Importantly, this assumption does not rule out that the effect might be zero. In our empirical application, it imposes that the effect of the true crop productivity on the outcome variable is non-negative. This assumption is motivated by the empirical literature because we assume that the articles using one proxy — the FAO-GAEZ version 3 — are at least identifying the correct sign of the true effect. If a researcher is not willing to impose this type of assumption, Appendix D proposes an alternative way to partially identify this coefficient without making assumptions about its sign.
@@
Assumption 4 imposes that the measurement error terms in our proxies of the true variable of interest are uncorrelated with the unobservable economic shock after controlling for observable covariates. [^8] This assumptions is implied by $Cov\left(A,\epsilon\right)=0$ for any $A\in\left\{W_{1},\ldots,W_{J},X^{*},Z_{1},Z_{2}\right\}$ as discussed in Appendix H. This assumption weakens the implicit identifying assumption in the empirical literature about agricultural productivity (Nunn and Qian, 2011; Bustos et al., 2016; Acharya et al., 2016) since applied researchers impose that $U_{1}=U_{2}=0$.
In our empirical application, Assumption 4 intuitively imposes that prediction errors in the FAO-GAEZ measures are uncorrelated with the unobserved economic forces that determine the outcome variable. This assumption would be violated if the institutional factors that impact $\epsilon_{res}$ positively also affect the quality of the environmental data that is used when constructing the FAO-GAEZ measure. In this case, regions with large values of $\epsilon_{res}$ would have a small amount of measurement error. Consequently, it is implausible that $Cov\left(U_{1},\epsilon_{res}\right)=0$.
@@
To illustrate the use of our partial identification strategy, we reevaluate the findings of Nunn and Qian (2011), Bustos et al. (2016), and Acharya et al. (2016). We find that the estimated effect of agricultural productivity on population may be substantially smaller than reported by Nunn and Qian (2011); that there may be no effect on political behavior, in contrast to the findings of Acharya et al. (2016); and that the effect on structural transformation remains robust, consistent with Bustos et al. (2016).
@@
Nunn and Qian (2011) study the impact of the introduction of the potato in the Old World on population. The empirical specification uses panel data to estimate Equation (8):
@@
Column (6) shows the results associated with our method. As we develop a partial identification strategy, our result is an estimated interval. This interval shows a wide range of possible effects, with an increase of one standard deviation in $\log(1+A_{j}^{potato})$ leading to an increase in population between 2.9% and 22.9%. Importantly, this interval exhausts all the information contained in the first and second moments of the data-generating process, implying that the large estimate for the upper bound is not a result of our estimator being loose. Such a wide estimated interval, especially in comparison to the small standard errors reported in Columns (4) and (5), suggests that model uncertainty with respect to the existence of non-classical measurement error is more concerning than sample uncertainty.
@@
Column (3) shows the results associated with our method. The estimated interval accounts for nonclassical measurement error in the explanatory variable and shows a wide range of possible effects. Interestingly, the estimated lower bound is almost a third of the point estimate associated with FAO-GAEZ version 3 (Column (1)) and is even lower than the point estimate associated with FAO-GAEZ version 4 (Column (4)), suggesting that accounting for measurement error in this empirical context is relevant. Lastly, when we account for sample uncertainty when estimating our bounds in Column (3), we do not reject the null of no effect at the 10%-confidence level. [^14] The estimated 90%-confidence interval is $\left[-0.006,0.257\right]$ while the 99%-confidence interval is $\left[-0.013,0.261\right]$. In Column (3) of Table 6, we report the intersection of these estimated intervals with $\mathbb{R}_{+}$ due to Assumption 3.
Recommendations to the Practitioner
In this section, we provide a detailed, 5-step guide on how to use our method in any empirical application with a mismeasured treatment variable and two proxies.
First, the practitioner must reflect on the exogeneity of the correctly measured treatment variable (Assumption 1). If the true treatment were observable, could the practitioner identify the treatment effect of interest using a linear regression model? This type of argument is present in any empirical application and practitioners still need to think about it when using our proposed method. If the answer is positive, proceed to the next step. If the answer is negative, the practitioner should use a different approach (e.g., instrumental variables).
Second, the practitioner must reflect on the exogeneity of the measurement errors (Assumption 4). Are the measurement error terms uncorrelated with unobserved factors in the outcome equation? To answer this question, it is important to think about the causes of the measurement error. Is it due to arbitrary modeling decisions about imputation or interpolation methods? Or is it due to better institutions being able to collect better data? If the first scenario is more likely, then Assumption 4 is plausible. If the practitioner believes that Assumption 4 is valid, proceed to the next step.
Third, the practitioner must reflect on their priors about the sign of the coefficient of interest (Assumption 3) and about the sign of the correlation between the measurement error terms (Assumption 5).
To think about the sign of the coefficient of interest, the practitioner may compare different economic theories or analyze previous empirical results using different datasets. Do they agree about the direction of the true effect? If so, Assumption 3 is plausible.
To think about the sign of the correlation between the measurement error terms, the practitioner must, once more, think about the causes of the measurement error. If it is hard to collect high-quality data on some regions, then Assumption 5 is plausible. If measurement error is due to different modeling choices, then Assumption 5 may not be plausible.
  1. 1.
    If the practitioner believes that Assumptions 3 and 5 are plausible, then proceed with the method described in Section 3.
  2. 2.
    If the practitioner believes that only Assumption 3 is plausible, then proceed with the method described in Appendix C.
  3. 3.
    If the practitioner believes that only Assumption 5 is plausible, then proceed with the method described in Appendix D.
Fourth, the practitioner must estimate the bounds around the true coefficient of interest using the estimator proposed in Appendix E.1. The estimated bounds are interesting for at least three aspects:
  1. 1.
    If all three terms in the estimated lower bound are similar to each other, then the results are robust to measurement error, and the practitioner may focus their analysis on their OLS results. This scenario is illustrated by our analysis of the dataset used by Bustos et al. (2016).
  2. 2.
    If the estimated lower bound is numerically close to zero, then the true coefficient of interest may be small. This type of result suggests that the economic relevance of the target parameter may not be large. This scenario is illustrated by our analysis of the dataset used by Acharya et al. (2016).
  3. 3.
    If the estimated interval is wide, then the model uncertainty generated by the possible existence of measurement error is relevant. This scenario is illustrated by our analysis of the dataset used by Nunn and Qian (2011).
Lastly, the practitioner must estimate the confidence intervals around the true coefficient of interest using the algorithm proposed in Appendix E.2. If the confidence interval intersects zero, then the “null hypothesis that the true coefficient of interest equals zero” is not rejected. In this case, the data does not provide evidence that there is an effect of $X^{*}$ on $Y$ when we account for non-classical measurement error.